OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). A)parallel: y = -x - 1 perpendicular: y = x + 2 B)parallel: y = x - 1 perpendicular: y = -x + 1 C)parallel: y = x + 1 perpendicular: y = -x - 3 D)parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

A)parallel: y = -x - 1 perpendicular: y = x + 2 B)parallel: y = x - 1 perpendicular: y = -x + 1 C)parallel: y = x + 1 perpendicular: y = -x - 3 D)parallel: y = 2x - 2 perpendicular: y = -2x - 1

Parth (parthkohli):

Parallel lines have the same slope, so you can eliminate two options.

OpenStudy (anonymous):

does that mean id eliminate d and a?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

ok now how would i find the right one of b and c?

Parth (parthkohli):

Perpendicular lines have the negative reciprocal. Which is -1.

Parth (parthkohli):

Actually you still have 2 choices.

OpenStudy (anonymous):

yeah cuz both the two choices left have a negative slope for the perpendicular

Parth (parthkohli):

Now you have to check if an equation passes through the point (-2,-1).

OpenStudy (anonymous):

does this mean my answer is C?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

Alright, the parallel has same slope while the perpendicular have opposite number and opposite sign of slope. Use this formula y - y1 = m(x - x1) y -(-1) = 1(x-(-2)) y=x+3 (parallel) y-y1 = m(x-x1) y-(-1) = -1(x-(-2)) y=-x-1

OpenStudy (anonymous):

i do not have an answer choice for that one tho, what do i do?

Parth (parthkohli):

Correct @mumina1999

OpenStudy (anonymous):

ok thank you , can i ask one more question?

OpenStudy (anonymous):

Ask two..

Parth (parthkohli):

Sure @mumina1999

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

Parth (parthkohli):

Ah, the point slope form would work.

Parth (parthkohli):

Can you calculate the slope?

OpenStudy (anonymous):

Two point form..

OpenStudy (anonymous):

let me try, i think point slope is this - y-y1 = m(x-x1)?

Parth (parthkohli):

Yes. Can you calculate the slope?

OpenStudy (anonymous):

yeah is it -9/3?

OpenStudy (anonymous):

Check it once again..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9/3 without the negative?

Parth (parthkohli):

\[{y_2 - y_1 \over x_2 - x_1}\]Just for a quick reminder.

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

thats what i did, ok now wat do i do

OpenStudy (anonymous):

Use the equation you have written above..

Parth (parthkohli):

Now, pick any point.

Parth (parthkohli):

You have two points; pick any.

OpenStudy (anonymous):

Use that point here : \[y- y_1 = m(x - x_1)\] You know m = 9/3 = ??

OpenStudy (anonymous):

y = 3x + 7?

Parth (parthkohli):

Let me check.\[y -1 = 3(x +3 )\]

Parth (parthkohli):

That's a 2*

Parth (parthkohli):

Yes, you're correct.

OpenStudy (anonymous):

Is not that right??

OpenStudy (anonymous):

ok but which answe choice would i choose becuase the answer choices seem to be in a different form

OpenStudy (anonymous):

y = 3x + 7 Rearrange it to form ax + by = c form..

OpenStudy (anonymous):

Can you do it??

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

3x - y = -7?

OpenStudy (anonymous):

Yep..

Parth (parthkohli):

Sorry, heh

OpenStudy (anonymous):

o ok thank you both!!!

Parth (parthkohli):

yw :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Keep practicing more questions related to this topic..

OpenStudy (anonymous):

yeah i will

Parth (parthkohli):

I forgot to practice and you can see what a duffer I am :P

OpenStudy (anonymous):

But those questions must be different from each other.. Ha ha ha..